# Optical next generation reservoir computing

### v2

@@
Artificial neural networks with internal dynamics exhibit remarkable capability in processing information. Reservoir computing (RC) is a canonical example that features rich computing expressivity and compatibility with physical implementations for enhanced efficiency. Recently, a new RC paradigm known as next generation reservoir computing (NGRC) further improves expressivity but compromises its physical openness, posing challenges for neuromorphic realizations. Here we demonstrate optical NGRC with large-scale computations performed by light scattering through disordered media. In contrast to conventional optical RC implementations, we drive our optical reservoir directly with time-delayed inputs. Much like digital NGRC that relies on polynomial features of delayed inputs, our optical reservoir also implicitly generates these polynomial features for desired functionalities. By leveraging the domain knowledge of the reservoir inputs, we show that the optical NGRC not only predicts the short-term dynamics of the low-dimensional Lorenz63 and large-scale Kuramoto-Sivashinsky chaotic time series, but also replicates their long-term ergodic properties. Optical NGRC shows superiority in shorter training length, increased interpretability and fewer hyperparameters compared to conventional optical RC, while achieving state-of-the-art forecasting performance. Given its scalability and versatility, our optical NGRC framework also paves the way for other next generation physical reservoirs, new applications and architectures in a broad sense.
@@ Introduction
Dynamical systems, which receive external stimuli and responsively react to them, possess remarkable capacity to manipulate and process information [1, 2, 3, 4]. As an artificial nonlinear dynamical system, reservoir computing (RC) is a type of recurrent neural networks (RNN) that makes ample use of its internal states by a weighted summation to achieve desired functionalities [5, 6, 7, 8, 9]. Owing to its hierarchical similarities to biological brains [10, 11, 12], along with the abilities to learn, adapt and memorize, RC has been widely recognized as a brain-like computing framework. It not only enables various applications in time series forecasting [7, 13, 14, 15], classification [16], prediction [17], attractor manipulation [18] and robots control [19], but also connects computing theory, machine learning, neuroscience, biology and physics broadly [20].
What makes RC so appealing is, in part, its physical compatibility. A broad range of physical mechanisms and substrates have been harnessed to implement reservoirs (Fig. 1a) [21], including analog electronics [22, 23], spintronic oscillators [24, 25], biological organoids [26], and many more. All of these physical implementations belong to neuromorphic computing, aiming for energy-efficient and high-throughput non-von Neumann architectures [27, 28, 29, 30]. Among others, optical computing is of particular interest [31, 32, 33], which employs photons as information carrier and light-matter interactions as processors, thereby exploiting the parallelism, energy efficiency and fast dynamics of light [34]. Within optical computing [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47], optical RC has a history of exploration for over a decade [48] and can be broadly classified into two types, i.e., delay-based reservoirs [49, 50, 51, 52, 53, 54, 55, 56] and spatial-distributed reservoirs [57, 58, 59, 60, 61, 62, 63, 64]. The former relies on either a single [49, 50, 51, 52, 53, 54, 55] or multiple [56] nonlinear devices with time-delayed feedback to create virtual reservoir nodes in the time domain. The latter encompasses versatile real reservoir systems built on semiconductor optical amplifiers [57], integrated delay line networks [58], diffractive optical elements [59], spatial light modulators (SLM) and cameras [60, 61], as well as multiple light scattering media [62, 63, 64].
@@
In this work, we demonstrate an optical NGRC scheme based on light scattering through disordered media. Specifically, we drive our optical system with time-delayed inputs (Fig. 1c), as opposed to feeding current inputs and reservoir states in almost all previous physical RC implementations. Instead of generating polynomial features directly as in digital NGRC, such a refinement also allows the optical setup to produce expanded polynomial features, embedded in the generic high-dimensional speckle intensity representations (Fig. 1e). Optical NGRC features a multitude of advantages over conventional counterparts. First, we demonstrate its efficacy in the short-term prediction of low-dimensional Lorenz63 and large-scale Kuramoto-Sivashinsky (KS) chaotic time series, achieving better prediction performance while using less than one tenth of the training data compared to previous state-of-the-art in optical RC [63]. Moreover, the optical NGRC also replicates the ‘climate’ of the original manifolds in the long-term dynamics, which acts as a photonic surrogate model. Furthermore, we show that the optical NGRC can accurately infer unmeasured state variables in observer prediction applications, outperforming standard digital interpolation methods. The optical NGRC demonstrated in this work delivers a clear interpretation of the neural network, that is, leveraging linear combinations of polynomial features from delayed inputs to empower versatile functionalities. Ultimately, though our scheme is an indirect form of digital NGRC, it offers substantial compatibility with physical computing systems, thereby paving the ways to tailor various other physical reservoirs.
@@ Results
where $\boldsymbol{W}_{in}$ is the input matrix mapping input data to the neuron domain, $\boldsymbol{W}_{r}$ is the interconnection matrix between neurons, $\boldsymbol{b}$ is the bias vector, and $f$ is the activation function that is typically nonlinear. To further control the memory of RC, many architectures also incorporate an additional hyperparameter, know as the leaky rate, to balance the current nonlinear activation with the previous reservoir state. After evolving the reservoir for a sufficient time, a linear estimator can be trained based on the reservoir states to produce the output $\boldsymbol{o}_{t}$ by $\boldsymbol{\hat{o}}_{t}\triangleq\boldsymbol{W}_{out}\boldsymbol{r}_{t}\approx\boldsymbol{o}_{t}$, where $\boldsymbol{W}_{out}$ is a readout layer mostly optimized through analytic linear regression (see Methods). Note that in the time series prediction tasks, the desired output is often the input, i.e., $\boldsymbol{o}_{t}=\boldsymbol{u}_{t}$. After training, the reservoir can autonomously evolve along a trajectory by closing the feedback loop in forecasting tasks. Importantly, the fixed nature of $\boldsymbol{W}_{in}$ and $\boldsymbol{W}_{r}$ renders RC a hardware-agnostic computing framework. RC is biologically plausible as it only trains the readout matrix $\boldsymbol{W}_{out}$, bypassing the challenges encountered in previous RNN training algorithms like backpropagation through time [1].
@@
where $1$ denotes the bias term, $\boldsymbol{u}_{t-k}\in\mathbb{R}^{M}$ is a delayed input from $k$ previous time steps ($k=1$ is used hereafter unless otherwise specified). $\otimes$ denotes the outer product and $\mathbb{U}$ is defined as an operation to collect all unique monomials from the matrix vectorization of the outer product.
@@
Figure 1: Optical next generation reservoir computing. a Reservoir computing (RC) is a special type of recurrent neural network and is compatible with various physical implementations. In the training phase, RC sequentially maps the current input ($\boldsymbol{u}_{t}$, blue) and the current reservoir state ($\boldsymbol{r}_{t}$) into the next reservoir state ($\boldsymbol{r}_{t+1}$, orange). After that, only a linear readout layer $\boldsymbol{W}_{out}$ is trained to match $\boldsymbol{\hat{o}}_{t+1}\triangleq\boldsymbol{W}_{out}\boldsymbol{r}_{t+1}$ with the desired output ($\boldsymbol{o}_{t+1}$, purple), which is often the input (i.e. $\boldsymbol{o}_{t+1}=\boldsymbol{u}_{t+1}$) in the time series prediction tasks. In the prediction phase, by feeding back the predicted output to the input, RC can autonomously evolve as a dynamical system. b Different from the conventional RC scheme, next generation reservoir computing (NGRC) directly synthesizes reservoir features by constructing the polynomial functions of the time-delayed inputs (e.g., $\boldsymbol{u}_{t}$ and $\boldsymbol{u}_{t-1}$), without relying on an actual reservoir. c Similar to the digital NGRC, the proposed optical NGRC also drives the optical reservoir with time-delayed inputs. The optical process generates the polynomial features of the inputs implicitly. d The schematic experimental setup for the optical NGRC. First, input data at the current and the previous time steps ($\boldsymbol{u}_{t}$ and $\boldsymbol{u}_{t-1}$) as well as a bias $\boldsymbol{b}$, are encoded onto the phase front of a laser beam via a spatial light modulator (SLM). Then, the modulated coherent light illuminates a disordered scattering medium, which provides rich mixing of the input and generates speckle patterns at the output. Finally, the reservoir features are obtained by measuring the intensity of the speckles in a camera. A computer (PC) is used to interface the SLM and the camera, as well as training and implementing readout layer. e The mathematical model of optical NGRC. The nonlinear, implicit reservoir speckle features $\boldsymbol{r}_{t+1}$ can be approximated as the multiplication of a system-given matrix $\boldsymbol{M}_{s}$ by a library of explicit polynomial feature terms $\boldsymbol{\Theta(U)}$.
@@
Figure 2: Optical NGRC for Lorenz63 attractor forecasting. a Time series of the Lorenz63 attractor (state variables $u_{1},u_{2},u_{3}$) that drives the optical NGRC. At each time step of the training phase, the input states from the current ($\boldsymbol{u}_{t}$) and the previous ($\boldsymbol{u}_{t-1}$) time steps are encoded to the optical system to generate reservoir features ($\boldsymbol{r}_{t+1}$). b The temporal evolution of 10 randomly selected optical reservoir nodes (out of 2,000 nodes), which resembles the dynamics of the input data. After training iterations of 4,000 time steps, a linear estimator $\boldsymbol{W}_{out}$ is trained to match the weighted sums of the reservoir features ($\boldsymbol{\hat{o}}_{t+1}=\boldsymbol{W}_{out}\boldsymbol{r}_{t+1}$) with the input data at the next time step ($\boldsymbol{u}_{t+1}$), i.e., $\boldsymbol{\hat{o}}_{t+1}\approx\boldsymbol{u}_{t+1}$. c Once $\boldsymbol{W}_{out}$ is optimized, the optical NGRC is switched to the autonomous mode and experimentally predicts short-term results for 400 time steps. The normalized root mean square error (NRMSE) over the first 5 time units of the prediction phase is 0.0971. d The optical NGRC projects onto a Lorenz63 attactor, experimentally obtained by the long-term forecasting results of 8,000 time steps. e The return map of the ground truth (blue) and the experimental prediction (red).
@@
With these in mind, we now construct optical NGRC as schematically shown in Figs. 1c-e. Our computing engine employs a continuous-wave laser as the light source, a phase-only SLM for data encoding, a scattering medium for information mixing, and a camera for feature detection (see Methods and Supplementary Note 1). Here, the input data from different time steps is encoded onto the spatial phase profile of light via the SLM. The scattering medium linearly connects the input and output optical fields via a transmission matrix, mixing the input as speckle patterns at the camera plane. Then, the formation of speckle feature vectors is analogous to analog random projection, which is a ubiquitous computation tool widely used in mathematics and signal processing [33]. Taking into account the nonlinear responses of phase encoding of the SLM ($x\rightarrow{\mathrm{exp}(ix)}$) and square-law detection of the camera ($x\rightarrow|x|^{2}$), the overall optical process defines the nonlinear mapping between the inputs ($\boldsymbol{u}_{t}$ and $\boldsymbol{u}_{t-1}$) and the reservoir state ($\boldsymbol{r}_{t+1}$) as:
@@
where $\boldsymbol{W}_{in1}$ and $\boldsymbol{W}_{in2}$ are random complex matrices given by the optical scattering medium. In contrast to conventional optical RC schemes where the reservoir state at the time step $t+1$ is calculated based on the current input $\boldsymbol{u}_{t}$ and the reservoir state $\boldsymbol{r}_{t}$ [49, 50, 51, 52, 53, 54, 55, 61, 62, 63], we replace $\boldsymbol{r}_{t}$ with the delayed input $\boldsymbol{u}_{t-1}$ (Fig. 1d). This modification generates implicitly the polynomial forms of input variables at time steps $t$ and $t-1$ (Fig. 1e), as evident by expanding $\boldsymbol{r}_{t+1}$ via Taylor series decomposition (see Supplementary Note 1):
@@
where $\boldsymbol{M}_{s}$ is a matrix given by the optical system, which mixes the underlying polynomial terms ($\boldsymbol{\Theta(U)}$) embedded in the speckle feature vector. In essence, the speckle vector can be understood as weighted sums of linear, quadratic and higher-order polynomial terms of $\boldsymbol{u}_{t}$ and $\boldsymbol{u}_{t-1}$. Stated differently, our optical system can compute similar feature terms just as the NGRC does in Eq. (2), only that an additional matrix linearly couples all these explicit terms together. Besides, the optimized linear readout matrix $\boldsymbol{W}_{out}$ trained in optical NGRC can be related to $\boldsymbol{W}_{out}^{\prime}$ in digital NGRC by $\boldsymbol{W}_{out}^{\prime}\approx\boldsymbol{W}_{out}\boldsymbol{M}_{s}$, thus validating that our optical implementation is equivalent to the digital NGRC operation.
@@
Forecasting Lorenz attractor. To demonstrate the effectiveness of the proposed optical NGRC, we firstly apply our setup to the low-dimensional Lorenz63 time series forecasting task (see Methods for dataset information). As illustrated in Fig. 2, we initially drive the optical system by encoding $[\boldsymbol{u}_{t},\boldsymbol{u}_{t-1}]^{T}$ onto the SLM and we gather in total 4,000 reservoir speckle feature vectors used for training (see Methods for experimental details). Figure 2b showcases the dynamics of 10 reservoir neurons measured in the experiment, providing nonlinear representations that reflect the characteristics of the input dataset. The smoothness of the reservoir dynamics, essential for reliable RC training [4], is guaranteed by the high stability of our experimental setup (see Supplementary Note 2). Then, we regress a digital readout layer $\boldsymbol{W}_{out}$ to map the reservoir state $\boldsymbol{r}_{t+1}$ to the next time step in the Lorenz63 attractor, i.e., $\boldsymbol{\hat{o}}_{t+1}\triangleq\boldsymbol{W}_{out}\boldsymbol{r}_{t+1}\approx\boldsymbol{u}_{t+1}$ (see Methods for training details). After $\boldsymbol{W}_{out}$ is obtained, the optical NGRC is used as an autonomous dynamical system for predicting another 400 time steps (see Supplementary Algorithm 1). The training and prediction are implemented in the optical experiment with an effective system frame rate of around 10 Hz (see Supplementary Note 2).
@@
In the short term, the optical NGRC shows decent forecasting capability of the Lorenz63 time series up to around 4 time units (Fig. 2c). Note that due to the nature of the chaotic systems, the prediction by optical NGRC would eventually diverge after a certain period of time, just as all models predicting chaos. Such a divergence does not imply the collapse of the RC model, rather, the ergodic (statistical) properties of the attractor are still preserved by RC, known as ‘climate’ replication [72, 73]. To this end, we run the trained optical NGRC for an extended period of 8,000 time steps. The long-term prediction consistently reproduces the manifold, as evident by the phase-space trajectory with double wings shown in Fig. 2d. Beyond visual inspection, we quantitatively evaluate the long-term forecasting performance by calculating the return map, in which the successive maxima of the third dimension $\boldsymbol{u}_{3}$ in time are collected and plotted. As shown in Fig. 2e, the experimentally obtained data points collectively cluster around the ground truth curve, albeit with a deviation due to the presence of the experimental noise and quantization.
@@
Figure 3: Optical NGRC for Kuramoto-Sivashinsky time series forecasting. a Experimental short-term prediction results of the Kuramoto-Sivashinsky (KS) time series with a domain size of $L=22$ and a spatial sampling of $S=64$. An optical NGRC with 2,500 optical reservoir nodes is used for KS forecasting, which employs the current ($\boldsymbol{u}_{t}$) and the previous ($\boldsymbol{u}_{t-1}$) time steps in each training iteration for a total training length of 6,000 time steps. The error subfigure (bottom) is the element-wise difference between the ground truth (top) and the experimental prediction (middle). The temporal axis is normalized by its largest Lyapunov time ($\lambda_{max}=0.043$). b A part of the long-term prediction results by optical NGRC (between $t_{1}$ and $t_{2}$, where the prediction starts at $t_{0}$). Albeit the complete deviation between the KS ground truth (top) and the optical NGRC predicted output (bottom) at the element-wise level, the optical NGRC replicates the long-term behavior of the KS chaotic system. c The power spectra of the long-term prediction in b (red), the KS ground truth (blue) and a random noise signal (yellow). The power spectra of the ground truth and optical NGRC predictions are in good agreement, in stark contrast to the power spectrum of the random noise background.
@@
Forecasting Kuramoto-Sivashinsky time series. Next, we use the optical NGRC in a more challenging scenario by forecasting another standard benchmark dataset in RC, i.e., a large-scale spatiotemporal chaotic KS time series (see Methods for information). In Fig. 3a, we illustrate the short-term prediction results obtained in the experiment through online Bayesian optimization (see Methods and Supplementary Note 2). After training the optical NGRC, the optical system can forecast the KS system reasonably well up to around 4 Lyapunov times (see definition in Methods), longer than the 2.5 Lyapunov times achieved previously [63]. At the prediction phase, the normalized root mean square (NRMSE) over the test period (6.45 Lyapunov times) is calculated as 0.2988 (see definition in Methods). We remark that, conventional RC typically necessitates a quite long warm-up period ranging from $100$ to $10,000$ time steps [73, 74], which can be challenging in situations where training data is limited or the physical evolution of the reservoir system is time-consuming. Thanks to the NGRC operation [70], we use in this study only 2 time steps for warm up and 6,000 time steps for training, much shorter than the total 90,500 time steps used in ref. [63]. Taken together, the better prediction performance in optical NGRC, combined with much less warm up and training data as well as a smaller reservoir size, collectively suggest the superiority of the optical NGRC over the conventional optical RC. The performance improvement using optical NGRC is also confirmed by numerical simulations (see Supplementary Note 3).

### v3

@@
Artificial neural networks with internal dynamics exhibit remarkable capability in processing information. Reservoir computing (RC) is a canonical example that features rich computing expressivity and compatibility with physical implementations for enhanced efficiency. Recently, a new RC paradigm known as next generation reservoir computing (NGRC) further improves expressivity but compromises its physical openness, posing challenges for realizations in physical systems. Here we demonstrate optical NGRC with computations performed by light scattering through disordered media. In contrast to conventional optical RC implementations, we drive our optical reservoir directly with time-delayed inputs. Much like digital NGRC that relies on polynomial features of delayed inputs, our optical reservoir also implicitly generates these polynomial features for desired functionalities. By leveraging the domain knowledge of the reservoir inputs, we show that the optical NGRC not only predicts the short-term dynamics of the low-dimensional Lorenz63 and large-scale Kuramoto-Sivashinsky chaotic time series, but also replicates their long-term ergodic properties. Optical NGRC shows superiority in shorter training length, increased interpretability and fewer hyperparameters compared to conventional optical RC based on scattering media, while achieving better forecasting performance. Our optical NGRC framework may inspire the realization of NGRC in other physical RC systems, new applications beyond time-series processing, and the development of deep and parallel architectures broadly.
@@ Introduction
Dynamical systems, which receive external stimuli and responsively react to them, possess remarkable capacity to manipulate and process information [1, 2, 3, 4]. As a nonlinear dynamical system, reservoir computing (RC) often builds upon recurrent neural networks (RNNs), utilizing the dynamics of internal reservoir states by a weighted summation to achieve desired functionalities [5, 6, 7, 8, 9]. The concept of RC not only finds various applications in time series forecasting [7, 10, 11, 12], classification [13], prediction [14], attractor manipulation [15] and robots control [16], but also connects computing theory, machine learning, neuroscience, biology and physics broadly [17].
What makes RC so appealing is, in part, its physical compatibility. A broad range of physical mechanisms and substrates have been harnessed to implement reservoirs (Fig. 1a) [18], including analog electronics [19, 20], spintronic oscillators [21, 22], biological organoids [23], and many more. All of these physical implementations aim at energy-efficient and high-throughput non-von Neumann architectures [24, 25, 26, 27]. Among others, optical computing is of particular interest [28, 29, 30], which employs photons as information carrier and light-matter interactions as processors, thereby exploiting the parallelism, energy efficiency and fast dynamics of light [31]. Within optical computing, optical RC has a history of exploration for over a decade [32, 33, 34] and can be broadly classified into two types, i.e., delay-based reservoirs [33, 34, 35, 36, 37, 38, 39, 40] and spatial-distributed reservoirs [32, 41, 42, 43, 44, 45, 46, 47]. The former relies on either a single [33, 34, 35, 36, 37, 38, 39] or multiple [40] nonlinear devices with time-delayed feedback to create virtual reservoir nodes in the time domain. The latter encompasses reservoir systems built on semiconductor optical amplifiers [32], integrated delay line networks [41], diffractively coupled optical elements [42], spatial light modulators (SLM) and cameras [43, 44], as well as multiple light scattering media [45, 46, 47].
@@
In this work, we demonstrate an optical NGRC scheme based on light scattering through disordered media. Specifically, we drive our optical system with time-delayed inputs (Fig. 1c), as opposed to feeding current inputs and reservoir states in almost all previous physical RC implementations. Instead of generating polynomial features directly as in digital NGRC, such a refinement also allows the optical setup to produce expanded polynomial features, embedded in the generic high-dimensional speckle intensity representations (Fig. 1e). Optical NGRC features a multitude of advantages over conventional optical RC [46] in processing time-series data. First, we demonstrate its efficacy in the short-term prediction of low-dimensional Lorenz63 and large-scale Kuramoto-Sivashinsky (KS) chaotic time series, achieving better prediction performance while using less than one tenth of the training data and a smaller reservoir compared to the previous state-of-the-art in optical RC [46]. Moreover, the optical NGRC also replicates the ‘climate’ of the original manifolds in the long-term dynamics, which acts as a photonic surrogate model. Furthermore, we show that the optical NGRC can accurately infer unmeasured state variables in observer prediction applications, outperforming standard digital interpolation methods. The optical NGRC demonstrated in this work delivers interpretable results by synthesizing features of time-delayed inputs through the optical system and then linearly combine them for versatile functionalities. Though our scheme is an indirect form of digital NGRC, it offers substantial compatibility with physical computing systems, thereby providing insights for tailoring various other physical reservoirs.
@@ Results
where $\boldsymbol{W}_{in}$ is the input matrix mapping input data to the neuron domain, $\boldsymbol{W}_{r}$ is the interconnection matrix between neurons, $\boldsymbol{b}$ is the bias vector, and $f$ is the activation function that is typically nonlinear. To further control the memory of RC, many architectures also incorporate an additional hyperparameter, know as the leaking rate, to balance the current nonlinear activation with the previous reservoir state. After evolving the reservoir for a sufficient time based on Eq. (1), a linear estimator can be trained to map the obtained reservoir states to the target outputs $\boldsymbol{o}_{t}$ by defining $\boldsymbol{\hat{o}}_{t}=\boldsymbol{W}_{out}\boldsymbol{r}_{t}\approx\boldsymbol{o}_{t}$, where $\boldsymbol{W}_{out}$ is a readout layer mostly optimized through analytic linear regression (see Methods). Note that in the time series prediction tasks, the desired output is often the input, i.e., $\boldsymbol{o}_{t}=\boldsymbol{u}_{t}$. After training, the reservoir can autonomously evolve along a trajectory by closing the feedback loop in forecasting tasks. Importantly, the fixed nature of $\boldsymbol{W}_{in}$ and $\boldsymbol{W}_{r}$ renders RC a hardware-agnostic computing framework. RC also bypasses the challenges encountered in previous RNN training algorithms, such as backpropagation through time [1], as it only trains the readout matrix $\boldsymbol{W}_{out}$.
@@
where $1$ denotes the bias term, $\boldsymbol{u}_{t-k}\in\mathbb{R}^{M}$ is a delayed input from $k$ previous time steps ($k=1$ is used hereafter unless otherwise specified). $\otimes$ denotes the outer product and $\mathbb{U}$ is defined as an operation to collect all unique monomials from the matrix vectorization of the outer product of two vectors.
@@
Figure 1: Optical next generation reservoir computing. (a) Reservoir computing (RC) is a special type of recurrent neural network and is compatible with various physical implementations. In the training phase, RC sequentially maps the current input ($\boldsymbol{u}_{t}$, blue) and the current reservoir state ($\boldsymbol{r}_{t}$) into the next reservoir state ($\boldsymbol{r}_{t+1}$, orange). After that, only a linear readout layer $\boldsymbol{W}_{out}$ is trained to match $\boldsymbol{\hat{o}}_{t}=\boldsymbol{W}_{out}\boldsymbol{r}_{t}$ with the desired output ($\boldsymbol{o}_{t}$, purple), which is often the input (i.e. $\boldsymbol{o}_{t}=\boldsymbol{u}_{t}$) in the time series prediction tasks. In the prediction phase, by feeding back the predicted output to the input, RC can autonomously evolve as a dynamical system. (b) Different from the conventional RC scheme, next generation reservoir computing (NGRC) directly synthesizes reservoir features by constructing the polynomial functions of the time-delayed inputs (e.g., $\boldsymbol{u}_{t}$ and $\boldsymbol{u}_{t-1}$), without relying on an actual reservoir. (c) Similar to the digital NGRC, the proposed optical NGRC also drives the optical reservoir with time-delayed inputs. The optical process generates the polynomial features of the inputs implicitly. (d) The schematic experimental setup for the optical NGRC. First, input data at the current and the previous time steps ($\boldsymbol{u}_{t}$ and $\boldsymbol{u}_{t-1}$) as well as a bias $\boldsymbol{b}$, are encoded onto the phase front of a laser beam via a spatial light modulator (SLM). Then, the modulated coherent light illuminates a disordered scattering medium, which provides rich mixing of the input and generates speckle patterns at the output. Finally, the reservoir features are obtained by measuring the intensity of the speckles in a camera. A computer (PC) is used to interface the SLM and the camera, as well as training and implementing readout layer. (e) The mathematical model of optical NGRC. The nonlinear, implicit reservoir speckle features $\boldsymbol{r}_{t+1}$ can be approximated as the multiplication of a system-given matrix $\boldsymbol{M}_{s}$ by a library of explicit polynomial feature terms $\boldsymbol{\Theta_{t}}$
@@
Figure 2: Optical NGRC for Lorenz63 attractor forecasting. (a) Time series of the Lorenz63 attractor (state variables $u_{1},u_{2},u_{3}$) that drives the optical NGRC. At each time step of the training phase, the input states from the current ($\boldsymbol{u}_{t}$) and the previous ($\boldsymbol{u}_{t-1}$) time steps are encoded to the optical system to generate reservoir features ($\boldsymbol{r}_{t+1}$). (b) The temporal evolution of 10 randomly selected optical reservoir nodes (out of 2,000 nodes), which resembles the dynamics of the input data. After training iterations of 4,000 time steps, a linear estimator $\boldsymbol{W}_{out}$ is trained to match the weighted sums of the reservoir features ($\boldsymbol{\hat{o}}_{t}=\boldsymbol{W}_{out}\boldsymbol{r}_{t}$) with the input data at the next time step ($\boldsymbol{u}_{t}$), i.e., $\boldsymbol{\hat{o}}_{t}\approx\boldsymbol{u}_{t}$. (c) Once $\boldsymbol{W}_{out}$ is optimized, the optical NGRC is switched to the autonomous mode and experimentally predicts short-term results for 400 time steps. The normalized root mean square error (NRMSE) over the first 5 time units of the prediction phase is 0.0971. (d) The optical NGRC projects onto an attractor similar to the Lorenz63 attactor, experimentally obtained by the long-term forecasting results of 8,000 time steps. (e) The return map of the ground truth (blue) and the experimental prediction (red).
@@
With these in mind, we now adapt NGRC into a format that is compatible with optical implementations (Figs. 1(c)-(e)), such that we can use a similar optical setup as for conventional optical RC [45, 46] for NGRC. Our computing engine employs a continuous-wave laser as the light source, a phase-only SLM for data encoding, a scattering medium for information mixing, and a camera for feature detection (see Methods and Supplementary Note 1). Here, the input data from different time steps is encoded onto the spatial phase profile of light via the SLM. The scattering medium linearly connects the input and output optical fields via a transmission matrix, mixing the input as speckle patterns at the camera plane. Then, the formation of speckle feature vectors is analogous to random projection, which is a ubiquitous computation tool widely used in mathematics and signal processing [30]. Taking into account the nonlinear responses of phase encoding of the SLM ($x\rightarrow{\mathrm{exp}(ix)}$) and square-law detection of the camera ($x\rightarrow|x|^{2}$), the overall optical process defines the nonlinear mapping between the inputs ($\boldsymbol{u}_{t}$ and $\boldsymbol{u}_{t-1}$) and the reservoir state ($\boldsymbol{r}_{t+1}$) as:
@@
where $\boldsymbol{W}_{in1}$ and $\boldsymbol{W}_{in2}$ are random complex matrices given by the optical scattering medium. In contrast to conventional optical RC schemes where the reservoir state at the time step $t+1$ is calculated based on the current input $\boldsymbol{u}_{t}$ and the reservoir state $\boldsymbol{r}_{t}$ [33, 34, 35, 36, 37, 38, 39, 44, 45, 46], we replace $\boldsymbol{r}_{t}$ with the delayed input $\boldsymbol{u}_{t-1}$ (Fig. 1(d)). Notably, this is different from the conventional RC framework augmented by delayed inputs [55, 56]. Such a modification generates implicitly the polynomial forms of input variables at time steps $t$ and $t-1$ (Fig. 1(e)), as evident by expanding $\boldsymbol{r}_{t+1}$ via Taylor series decomposition (see Supplementary Note 1):
@@
where $\boldsymbol{M}_{s}$ is a matrix given by the optical system, which mixes the underlying polynomial terms ($\boldsymbol{\Theta_{t}}$) embedded in the speckle feature vector. In essence, the speckle vector can be understood as weighted sums of linear, quadratic and higher-order polynomial terms of $\boldsymbol{u}_{t}$ and $\boldsymbol{u}_{t-1}$. Stated differently, our optical system can compute similar feature terms just as the NGRC does in Eq. (2), only that an additional matrix linearly couples all these explicit terms together. Besides, the optimized linear readout matrix $\boldsymbol{W}_{out}$ trained in optical NGRC can be related to $\boldsymbol{W}_{out}^{\prime}$ in digital NGRC by $\boldsymbol{W}_{out}^{\prime}\approx\boldsymbol{W}_{out}\boldsymbol{M}_{s}$, thus validating that our optical implementation is equivalent to the digital NGRC operation (see Supplementary Note 1). Due to the presence of the system-given matrix $\boldsymbol{M}_{s}$, the optical implementation operates in an indirect manner. This indirect way of implementing NGRC optically also inherits the advantage of interpretability of digital NGRC, as we can explain the reservoir computations via the synthesized features of the time-delayed inputs and then a readout layer $\boldsymbol{W}_{out}^{\prime}$ linearly combining them for specific tasks.
@@
Forecasting Lorenz attractor. To demonstrate the effectiveness of the proposed optical NGRC, we firstly apply our setup to the low-dimensional Lorenz63 time series forecasting task (see Methods for dataset information). As illustrated in Fig. 2, we initially drive the optical system by encoding $[\boldsymbol{u}_{t},\boldsymbol{u}_{t-1}]^{T}$ with a time interval of $\Delta t=0.025$ onto the SLM and we gather in total 4,000 reservoir speckle feature vectors used for training (see Methods for experimental details). Figure 2(b) showcases the dynamics of 10 reservoir neurons measured in the experiment, providing nonlinear representations that reflect the characteristics of the input dataset. The smoothness of the reservoir dynamics, essential for reliable RC training [4], is guaranteed by the high stability of our experimental setup (see Supplementary Note 2). Then, we regress a digital readout layer $\boldsymbol{W}_{out}$ to map the reservoir state $\boldsymbol{r}_{t}$ to the next time step in the Lorenz63 attractor, i.e., $\boldsymbol{\hat{o}}_{t}=\boldsymbol{W}_{out}\boldsymbol{r}_{t}\approx\boldsymbol{u}_{t}$ (see Methods for training details). After $\boldsymbol{W}_{out}$ is obtained, the optical NGRC is used as an autonomous dynamical system for predicting another 400 time steps (see Supplementary Algorithm 1). The training and prediction are implemented in the optical experiment with an effective system frame rate of around 10 Hz (see Supplementary Note 2).
@@
In the short term, the optical NGRC shows decent forecasting capability of the Lorenz63 time series up to around 4 time units (Fig. 2(c)). Note that due to the nature of the chaotic systems, the prediction by optical NGRC would eventually diverge after a certain period of time, just as all models predicting chaos. Such a divergence does not imply the collapse of the RC model, rather, the ergodic (statistical) properties of the attractor are still preserved by RC, known as ‘climate’ replication [57, 58]. To this end, we run the trained optical NGRC for an extended period of 8,000 time steps. The long-term prediction consistently reproduces the manifold, as evident by the phase-space trajectory with double wings shown in Fig. 2(d). Beyond visual inspection, we quantitatively evaluate the long-term forecasting performance by calculating the return map, in which the successive maxima of the third dimension $\boldsymbol{u}_{3}$ in time are collected and plotted. As shown in Fig. 2(e), the experimentally obtained data points collectively cluster around the ground truth curve, albeit with a deviation due to the presence of the experimental noise and quantization.
@@
Figure 3: Optical NGRC for Kuramoto-Sivashinsky time series forecasting. (a) Experimental short-term prediction results of the Kuramoto-Sivashinsky (KS) time series with a domain size of $L=22$ and a spatial sampling of $S=64$. An optical NGRC with 2,500 optical reservoir nodes is used for KS forecasting, which employs the current ($\boldsymbol{u}_{t}$) and the previous ($\boldsymbol{u}_{t-1}$) time steps in each training iteration for a total training length of 6,000 time steps. The error subfigure (bottom) is the element-wise difference between the ground truth (top) and the experimental prediction (middle). The temporal axis is normalized by its largest Lyapunov time ($\lambda_{max}=0.043$). (b) A part of the long-term prediction results by optical NGRC (between $t_{1}$ and $t_{2}$, where the prediction starts at $t_{0}$). Albeit the complete deviation between the KS ground truth (top) and the optical NGRC predicted output (bottom) at the element-wise level, the optical NGRC replicates the long-term behavior of the KS chaotic system. (c) The power spectra of the long-term prediction in (b) (red), the KS ground truth (blue) and a random noise signal (yellow). The power spectra of the ground truth and optical NGRC predictions are in good agreement, in stark contrast to the power spectrum of the random noise background.
@@
Forecasting Kuramoto-Sivashinsky time series. Next, we use the optical NGRC in a more challenging scenario by forecasting another standard benchmark dataset in RC, i.e., a large-scale spatiotemporal chaotic KS time series (see Methods for dataset information). In Fig. 3(a), we illustrate the short-term prediction results obtained in the experiment through online Bayesian optimization (see Methods and Supplementary Note 2). After training the optical NGRC, the optical system can forecast the KS system reasonably well up to around 4 Lyapunov times (see definition in Methods), longer than the 2.5 Lyapunov times achieved previously [46]. At the prediction phase, the normalized root mean square (NRMSE) over the test period (6.45 Lyapunov times) is calculated as 0.2988 (see definition in Methods). We remark that, conventional RC typically necessitates a quite long warm-up period ranging from $100$ to $100,000$ time steps [58, 59], which can be challenging in situations where training data is limited or the physical evolution of the reservoir system is time-consuming. Thanks to the NGRC operation [53], we use in this study only 2 time steps for warm up and 6,000 time steps for training, much shorter than the total 90,500 time steps used in ref. [46]. Taken together, the better prediction performance in optical NGRC, combined with much less warm up and training data as well as a smaller reservoir size, collectively suggest the superiority of the optical NGRC over the conventional optical RC based on scattering media. The performance improvement using optical NGRC is also confirmed by numerical simulations (see Supplementary Note 3).